24. The demand curve and supply curve for one-year discount

bonds with a face value of $1,000 are represented by the

following equations:

Bd : Price = -0.8 * Quantity + 1100

Bs : Price = Quantity + 680

a. What is the expected equilibrium price and quantity

of bonds in this market?

b. Given your answer to part (a), what is the expected

interest rate in this market?


1
Expert's answer
2022-06-13T08:21:53-0400

"\begin{cases}
 P=-0.8Q+1100, \\
 P=Q+680;
\end{cases}"

"-0.8Q+1100=Q+680,"

"1.8Q=420,"

"Q=233,"

"P=913."